If you decrease the distance between two celestial objects by a factor of 3.5, how does the resultant gravitational force between the two objects compare to the original force? The force is ____ times stronger.
Answer #1
  • The force is __12.25__ times stronger.

Gravitational force (F)= km1m2/r^2

So F∝1/x^2

F1×x1^2 =F2×x2^2

F1×x1^2 =F2×(x1/3.5)^2

F2=12.25F1.